OpenStudy (anonymous):

Use the First Fundamental Theorem of Calculus derivative of 3sqrt1-s^2 ds from 0 to sinx

OpenStudy (anonymous):

\[\int\limits_{0}^{sinx} \sqrt[3]{1-s ^{2}} ds\]

OpenStudy (anonymous):

So any one???? pls

OpenStudy (anonymous):

\[ \cos(x) \sqrt[3]{1-\sin(x) ^{2}} \]\

ganeshie8 (ganeshie8):

\(\large F(\sin x) = \frac{d}{dx} \int_0^{\sin x}\sqrt[3]{1-x^2} ds \)